1. In the contest of Serach for Most Outstanding Teacher, 3 candidates are teaching Mathematics, 2 are teaching English, 2 are teaching Science and 1 is teaching in Filipino. What is the probability that the winner is teaching:
a. Mathematibs
b. Science
c. Filipino
d. Any subject except English
e. any subject except math
2. If the probabilities are 0.40, 0.55 and 0.3 that student will get a failing grade either in Accountancy or Marketing, or in both, what is the probability that he fail at least one of these subjects?

Answer by stanbon(48568)

1. In the contest of Serach for Most Outstanding Teacher, 3 candidates are teaching Mathematics, 2 are teaching English, 2 are teaching Science and 1 is teaching in Filipino. What is the probability that the winner is teaching:
a. P(Mathematics)= 3/8
b. P(Science)=2/8
c. P(Filipino)=1/8
d. P(Any subject except English)=1-(2/8)=6/8
e. P(any subject except math)=1-(3/8)=5/8
-----------------
2. If the probabilities are 0.40, 0.55 and 0.3 that student will get a failing grade either in Accountancy or Marketing, or in both, what is the probability that he fail at least one of these subjects?
P(fail at least one)=?
P(fail Account OR Market)= P(fail Acc)+P(fail Mar)-P(fail both)=0.4+0.55-0.3=0.65
P(fail at least one)=0.65
============
Cheers,
Stan H.